OpenStudy (anonymous):

The radius of a circle is multiplied by 7. Its area is _____ times bigger than the original circle. 7 21 49 70

OpenStudy (saifoo.khan):

7^2 = 49

OpenStudy (anonymous):

\[A = \pi r ^{2}\] So \[A _{1} = \pi r ^{2}\] \[A _{2} = \pi (7r)^{2}\] \[A _{2} / A _{1} = (\pi (7r)^{2}) / (\pi r ^{2}) = 49 (\pi r ^{2} / \pi r^{2}) = 49\]